# A Comprehensive Study on Large Language Models for Mutation Testing

### v4

@@ 3.2. RQ2: Performance on Validity
To evaluate the validity of mutation generation approaches, we consider the aspects that may impact their validity, i.e., compilability rate, duplicate mutation rate, and equivalent mutation rate. This RQ is divided into the following sub-RQs.
@@
The column Comp. Rate of Table 4 presents the overall Compilability Rate () of each approach. Among the traditional approaches, the rule-based approaches, PIT and Major, demonstrate superior compilability, with PIT achieving a perfect 100% , and Major following closely at 97.60%. Note that although Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true). PIT uses the mutation operators more restrictively and previously filters out non-compilable mutations. In contrast, LEAM and BERT, have rates at 35.0% and 22.5%, respectively. For LLM-based approaches, there is a considerable variation in compilability. The existing LLM-based approaches, BugFarm and LLMorpheus, achieve relatively high rates of 74.1% and 67.8%, respectively. Among the LLMs using our prompt, DeepSeek-V3-761b attains the highest at 77.6%, followed closely by GPT-4o (76.4%), DeepSeek-Coder-V2-236b (75.8%), and GPT-4o-Mini (75.3%). Compared to small models (i.e., LEAM and BERT), GPTs and DeepSeek models show significantly higher compilation success rates. For example, the compilability rate of GPT-4o is over 30% higher than that of LEAM and BERT.
3.2.2. RQ2.2: Duplicate Mutation Rate
@@
As aforementioned, duplicate mutations refer to those that are redundant or identical to the original code, providing no meaningful contribution to the mutation testing process. Therefore, a lower Duplicate Mutation Rate () is desirable, as it indicates a more efficient mutation generation process with minimal redundancy. The column Dup. Mut. Rate of Table 4 presents the results.
@@
Rule-based approaches, PIT and Major, achieve a perfect 0% . This is expected, as these approaches apply well-defined mutation rules that ensure each generated mutant introduces a meaningful change to the program. Other traditional approaches, i.e., LEAM and BERT, show relatively low s of 2.3% and 1.7%, respectively. In contrast, LLMs generate a significant portion of duplicate mutations, exhibiting a higher . For the LLMs from OpenAI, GPT-3.5, GPT-4o, and GPT-4o-Mini exhibit moderate s of 10.3%, 7.8%, and 7.0%, respectively. For the existing LLM-based approaches, BugFarm and LLMorpheus, exhibit similar s of 7.5% and 7.1%, respectively. The open-source LLMs such as StarCoder and DeepSeek-Coder-V2-236b demonstrate higher s of 12.8% and 7.7%, respectively, while CodeLlama shows the highest at 37.1%.
3.2.3. RQ2.3: Equivalent Mutations
@@
Given the undecidable nature of identifying equivalent mutations, we manually evaluate each mutation’s equivalence, following the methodology of the existing study (Kushigian et al., 2024). For each mutation generation approach, we randomly sample a subset of compilable mutations, ensuring a 95% confidence level with a 10% margin of error.
@@
Two authors independently evaluate the equivalence of the selected mutations to minimize subjectivity. The procedure for checking a mutation involves first transplanting it into the original project and executing all tests. If it passes all tests, the author then manually verifies whether the mutation is semantically equivalent. Note that PIT does not generate source-level mutations, so its mutations must be manually extracted from Java bytecode and transformed back into source-level representations, which is significantly more time-consuming.
@@
In total, 1247 mutations were labeled, 709 of which were filtered by tests (i.e., killed), and the remaining 538 were manually checked. The entire process took approximately twelve hours per person. The manual verification requires checking the code context, which may involve multiple classes and methods. For example, to check a mutation that replaces an integer with a method invocation, where the method is invoked by the this object but defined in the parent class, we should review the parent class file and verify whether its return value is consistently equal to the integer. We measure the consistency between the two authors by Cohen’s kappa coefficient () (Viera et al., 2005), achieving an overall value of 0.8, which indicates a high level of agreement. In cases of disagreement, both authors conducted thorough discussions to resolve contradictions and finally reach consistent labels. The labeling results are available at our repo.
@@
Table 10 presents the estimated Equivalent Mutation Rate (), reporting the number of sampled mutations (#Sampled), the number of equivalent mutations we identified (#Eq. Mut), and the for each approach. Among the traditional approaches, PIT achieves the lowest of 1.0%, followed by Major with an of 2.1%, indicating that rule-based techniques produce fewer equivalent mutations. In contrast, the LLM-based approaches exhibit higher s, with variations across different models. StarChat has the highest s of 10.6%, while GPT-4o, GPT-4o-Mini, and DeepSeek-Coder-V2-236b show relatively lower s, ranging from 3.1% to 5.2%. Overall, although the OpenAI models such as GPT-4o and GPT-4o-Mini tend to have slightly higher s than PIT and Major, the increase remains modest and within an acceptable range.
Note that the estimated of Major given by the recent study of Kushigian et al. (Kushigian et al., 2024) is 2.97%, where the authors sampled a subset of mutations for manual judgment. We adopt the same experimental process and the result of Major is 2.0%, which is consistent with theirs, considering the randomness introduced by sampling. Moreover, the definition of equivalent mutations in this paper is the set of syntactically different but semantically equivalent mutations, which may differ from existing studies (Tip et al., 2024).
@@ 3.3. RQ3: Performance on Efficiency
Average Generation Time: The column Time of Table 4 presents the results of Average Generation Time in seconds. Among the traditional appraoches, PIT demonstrates the fastest generation time at 0.02 seconds per mutation, followed by Major, which takes 0.08 seconds, making them the most efficient approaches. In contrast, LEAM and BERT take 3.06 seconds and 2.34 seconds, respectively. The existing LLM-based approaches, Among the LLMs, the LLMs of OpenAI (i.e., GPT-3.5, GPT-4o, and GPT-4o-Mini) show relatively faster generation, ranging from 1.31 to 1.79 seconds. The open-source LLMs StarChat, CodeLlama, and DeepSeek require 7.53, 9.06, and 4.25 seconds, respectively. The results suggest that the rule-based approaches are the most time-efficient for generating mutations, which are tens to hundreds of times faster than other approaches. Note that the time costs of self-deployed open-source models typically depend on the types of GPUs used for execution. Additionally, the response time of models accessed via network APIs may vary across service providers and can be further affected by network fluctuations. We further discuss these potential impacts in the Section 5.
| Approach | Avg. Input Token | Avg. Output Token |
| BugFarm (GPT-4o) | 200 | 368 |
| LLMorpheus (GPT-4o) | 370 | 74 |
| LLMut (GPT-3.5) | 126 | 59 |
| LLMut (GPT-4o) | 126 | 59 |
| LLMut (GPT-4o-M) | 126 | 59 |
| LLMut (DS-761b) | 126 | 58 |
| LLMut (DS-236b) | 126 | 59 |
| LLMut (SC-16b) | 126 | 63 |
| LLMut (CL-13b) | 126 | 58 |
Average Token: Table 11 presents the average number of input and output tokens used by each LLM-based approach to generate one mutation. The existing LLM-based approaches, BugFarm and LLMorpheus, consume 200 and 370 input tokens, and 368 and 74 output tokens per mutation, respectively. LLMs using our prompt consume substantially fewer tokens, averaging 126 input tokens and 59–63 output tokens per mutation. As all LLMs share identical inputs, their input token counts remain the same.
3.4. RQ4: Categorizations of Non-Compilable Mutations
To answer this RQ, we first identify the error types of non-compilable mutations and then analyze the surrounding code context types that are error-prone to generate them.
@@ 3.4.1. Categorizations of compilation error types
Non-compilable mutations require a compilation step to filter out, which results in wasted computational resources. As mentioned in Section 3.2, LLMs generate many non-compilable mutations. This RQ analyzes the types of errors and potential root causes of these non-compilable mutations. Following the setting of the previous steps, we first sample 384 non-compilation mutations from the outputs of GPT models, ensuring the confidence level is 95% and the margin of error is 5% (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015). From the manual analysis of these non-compilable mutations, we identified nine distinct error types, shown as follows.
@@
Incorrect Location: This type of error is due to a mutation being applied to an incorrect code location.
Incorrect Exceptions: This type of error is due to throwing illegal exceptions.
@@
3.4.2. Categorizations of code context of non-compilable mutations
@@
To analyze which types of code are prone to causing LLMs to generate non-compilable mutations, we examined the code locations of all non-compilable mutations generated by all nine mutation generation approaches in Table 4, as shown in Figure 3. For most approaches (except Major and BugFarm), the code locations with MemberReference and MethodInvocation are the top-2 most-prevent AST node types. For example, for GPT and DeepSeek models, about 40% of non-compilable mutations occur at the location with MemberReference or MethodInvocation. This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references. If any required method or member is missing or incorrectly specified, it can easily lead to non-compileable mutations. The errors highlight a need for better context-aware mutation generation, ensuring that method calls and member references align with the intended program structure. Additionally, we inspect the deletion mutations rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 3.4% (DeepSeek-V3-761b), 1.11% (DeepSeek-Coder-V2-236b), 3.4% (BugFarm), 0.8% (LLMorpheus), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% (BERT), and 16.61% (Major) of all their non-compilable mutations, respectively. Thus, for LLMs, deletion is not the major reason for non-compilation.
3.5. RQ5: Distributions of Alive Mutations
@@
As shown in Table 4, the mutation scores of rule-based approaches (i.e., PIT and Major) are lower than LLMs, indicating their mutations are harder to detect. Therefore, we intend to analyze the reasons for survival mutations. To address this RQ, we adopt a setting similar to the analysis of equivalent mutations. Specifically, for each approach, we randomly sample its alive mutations and manually classify them. We employ the categories of alive mutation causes defined in the study by Just et al. (Just et al., 2014a) as follows.
@@
(1) Not covered, i.e., the mutation is located in code never executed by the test suite. We refer to the mutation set of this category as .
@@
(2) Covered but no state infection, i.e., the mutation is executed but does not propagate any state change. We refer to the mutation set of this category as .
@@
(3) State infection without detection, i.e., the mutation alters the program state but any test oracle does not observe the deviation. We refer to the mutation set of this category as .
@@
These categories are mutually exclusive, with the following relationships: , , and . This classification enables us to analyze the underlying reasons for alive mutations and further evaluate the extent to which different approaches contribute to strengthening the test suite. Given an alive mutation, we first determine whether it is equivalent; if so, it is discarded. For non-equivalent mutations, we then check whether they are uncovered. If a mutation is covered, we further examine whether it alters the program state, i.e., whether any memory values are modified. Finally, the remaining alive mutations that are covered but do not lead to detectable differences are classified as not detected. This stepwise process ensures that every alive mutation is uniquely assigned to one category (equivalent, uncovered, not injecting state, or not detected).
@@
As shown in Table 13, the distributions of alive mutations differ significantly across approaches, where the columns NC%, NI%, and ND% represent the proportions of , , and within all non-equivalent mutations, respectively. First, LLM-based methods exhibit a substantially higher proportion of (i.e., not covered) mutations, meaning that more of their alive mutations reside in previously uncovered code lines. For instance, among the alive mutations generated by GPT-3.5 and BugFarm, 79.52% and 72.94% fall into the category, respectively, whereas PIT and Major exhibit much lower proportions of 46.24% and 45.65%. This characteristic highlights the potential of LLMs in guiding test suite augmentation by directing attention to untested program areas. Second, traditional approaches, especially rule-based methods such as PIT and Major, produce a larger fraction of (i.e., not injected) and (i.e., not detected) mutations compared to LLM-based methods. For instance, PIT and Major exhibit proportions of 29.03% and 32.61%, respectively, whereas all LLMs based on our prompt remain below 20%. These types of alive mutations require not only generating inputs that execute the mutated code but also propagating and detecting subtle semantic differences, which is inherently more challenging than increasing coverage. Therefore, the overall implication is that LLM-generated alive mutations are more directly beneficial for strengthening unit tests, as they emphasize extending coverage, while rule-based approaches tend to stress the harder task of state propagation and detection.

### v5

@@ 3.2. RQ2: Performance on Validity
To evaluate the validity of mutant generation approaches, we consider the aspects that may impact their validity, i.e., compilability rate, duplicate mutant rate, and equivalent mutant rate. This RQ is divided into the following sub-RQs.
@@
The column Comp. Rate of Table 4 presents the overall Compilability Rate () of each approach. Among the traditional approaches, the rule-based approaches, PIT and Major, demonstrate superior compilability, with PIT achieving a perfect 100% , and Major following closely at 97.60%. Note that although Major operates under simple syntax rules, it can still generate non-compilable mutants. For example, the Java compiler’s analyzer rejects constructs like while(true). PIT uses the mutation operators more restrictively and previously filters out non-compilable mutants. In contrast, LEAM and BERT, have rates at 35.0% and 22.5%, respectively.
For LLM-based approaches, there is a considerable variation in compilability. Among all LLM-based approaches, LLMut (DS-671b) and BugFarm (DS-671b) attain the highest of 77.6% and 77.5%, respectively. LLMut (GPT-4o), LLMut (DS-236b), and LLMut (GPT-4o-M) closely follow by achieving 76.4%, 75.8%, 75.3%, respectively. Most LLM-based approaches show significantly higher compilation success rates compared to the traditional learning-based approaches, such as LEAM and BERT. For example, the s of BugFarm, LLMorpheus, and LLMut with GPT and DeepSeek models are over 30% higher than those of LEAM and BERT.
3.2.2. RQ2.2: Duplicate Mutant Rate
@@
As aforementioned, duplicate mutants refer to those that are redundant or identical to the original code, providing no meaningful contribution to the mutant testing process. Therefore, a lower Duplicate Mutant Rate () is desirable, as it indicates a more efficient mutant generation process with minimal redundancy. The column Dup. Mut. Rate of Table 4 presents the results.
@@
Rule-based approaches, PIT and Major, achieve a perfect 0% . This is expected, as these approaches apply well-defined mutation rules that ensure each generated mutant introduces a meaningful change to the program. Other traditional approaches, i.e., LEAM and BERT, show relatively low s of 2.3% and 1.7%, respectively. In contrast, LLMs generate a significant portion of duplicate mutants, exhibiting a higher . For the LLM-based approaches evaluated with the best-performing model, DS-671b, the duplicate mutation rates (s) are as follows: BugFarm (DS-671b): 7.5%, LLMorpheus (DS-671b): 6.9%, and LLMut (DS-671b): 7.5%. The smaller models on LLMut exhibit much worse performance, such as LLMut (SC-16b) and LLMut (CL-13b). These results indicate that even equipped with the state-of-the-art LLM, these approaches still generate a non-negligible proportion of duplicate mutants.
3.2.3. RQ2.3: Equivalent Mutants
@@
Given the undecidable nature of identifying equivalent mutants, we manually evaluate each mutant’s equivalence, following the methodology of the existing study (Kushigian et al., 2024). For each mutant generation approach, we randomly sample a subset of compilable mutants, ensuring a 95% confidence level with a 10% margin of error.
@@
Two authors independently evaluate the equivalence of the selected mutants to minimize subjectivity. The procedure for checking a mutant involves first transplanting it into the original project and executing all tests. If it passes all tests, the author then manually verifies whether the mutant is semantically equivalent. Note that PIT does not generate source-level mutants, so its mutants must be manually extracted from Java bytecode and transformed back into source-level representations, which is significantly more time-consuming.
@@
In total, 1437 mutants were labeled, 860 of which were filtered by tests (i.e., killed), and the remaining 577 were manually checked. The entire process took approximately thirteen hours per person. The manual verification requires checking the code context, which may involve multiple classes and methods. For example, to check a mutant that replaces an integer with a method invocation, where the method is invoked by the this object but defined in the parent class, we should review the parent class file and verify whether its return value is consistently equal to the integer. We measure the consistency between the two authors by Cohen’s kappa coefficient () (Viera et al., 2005), achieving an overall value of 0.8, which indicates a high level of agreement. In cases of disagreement, both authors conducted thorough discussions to resolve contradictions and finally reach consistent labels. The labeling results are available at our repo.
@@
Table 10 presents the estimated Equivalent Mutant Rate (), reporting the number of sampled mutants (#Sampled), the number of equivalent mutants we identified (#Eq. Mut), and the for each approach. Among the traditional approaches, PIT achieves the lowest of 1.0%, followed by Major with an of 2.1%, indicating that rule-based techniques produce fewer equivalent mutants. In contrast, the LLM-based approaches exhibit higher s, with variations across different models. LLMut (SC-16) performs the worst by attaining the highest s of 10.6%, while LLMut (GPT-4o), LLMut (GPT-4o-M), and LLMut (DS-236b) show relatively lower s, ranging from 3.1% to 5.2%. BugFarm and LLMorpheus exhibit s ranging from 3.1% to 5.2%. Overall, the LLM-based approaches with advanced models (e.g., GPT-4o or DS-671b) tend to have slightly higher s than PIT and Major, the increase remains modest and within an acceptable range.
Note that the estimated of Major given by the recent study of Kushigian et al. (Kushigian et al., 2024) is 2.97%, where the authors sampled a subset of mutants for manual judgment. We adopt the same experimental process, and the result of Major is 2.0%, which is consistent with theirs, considering the randomness introduced by sampling. Moreover, the definition of equivalent mutants in this paper is the set of syntactically different but semantically equivalent mutants, which may differ from existing studies (Tip et al., 2025).
@@ 3.3. RQ3: Performance on Efficiency
Average Generation Time: The column Time of Table 4 presents the results of Average Generation Time in seconds. Among the traditional approaches, PIT demonstrates the fastest generation time at 0.02 seconds per mutant, followed by Major, which takes 0.08 seconds, making them the most efficient approaches. In contrast, LEAM and BERT take 3.06 seconds and 2.34 seconds, respectively. The LLM-based approaches present much slower generation speed, ranging from 1.31 to 9.06 seconds. In particular, all LLM-based approaches require more time when adopting open-source models. For example, LLMut (SC-16b), LLMut (CL-13b), and LLMut (DS-236b) require 7.53, 9.06, and 4.25 seconds, respectively. The results suggest that the rule-based approaches are the most time-efficient for generating mutants, which are tens to hundreds of times faster than other approaches. Note that the time costs of self-deployed open-source models typically depend on the types of GPUs used for execution. Additionally, the response time of models accessed via network APIs may vary across service providers and can be further affected by network fluctuations. We further discuss these potential impacts in Section 5.
| Approach | Avg. Input Token | Avg. Output Token | All Input Token | All Output Token |
| BugFarm (GPT-4o) | 200 | 368 | 2,780,213 | 4,807,979 |
| BugFarm (DS-671b) | 200 | 366 | 2,780,213 | 5,085,936 |
| LLMorpheus (GPT-4o) | 370 | 74 | 7,921,356 | 1,476,568 |
| LLMorpheus (DS-671b) | 370 | 73 | 7,921,356 | 1,558,988 |
| LLMut (GPT-3.5) | 126 | 59 | 4,353,426 | 1,855,042 |
| LLMut (GPT-4o) | 126 | 59 | 4,353,426 | 1,875,810 |
| LLMut (GPT-4o-M) | 126 | 59 | 4,353,426 | 2,027,850 |
| LLMut (DS-671b) | 126 | 58 | 4,353,426 | 1,947,224 |
| LLMut (DS-236b) | 126 | 59 | 4,353,426 | 2,064,658 |
| LLMut (SC-16b) | 126 | 63 | 4,353,426 | 1,766,537 |
| LLMut (CL-13b) | 126 | 58 | 4,353,426 | 1,599,127 |
Average Token: Table 11 presents the average number of input and output tokens used by each LLM-based approach to generate a single mutant. All the token usage is computed as the official library of OpenAI, Tiktoken666https://github.com/openai/tiktoken, which is the standard practice for estimating model usage.
From the table, we observe substantial differences across prompting strategies. BugFarm uses relatively short prompts (200 input tokens on average), but its output token usage is high (over 360 tokens) because it returns the code of the entire mutated method. Consequently, its overall token consumption needs over 2.7M input tokens and about 5M output tokens. In contrast, LLMorpheus requires the largest number of input tokens (370 on average, 7.9M in total), as its prompt includes up to 200 lines of surrounding context, yet it produces moderate output (around 74 tokens on average, 1.5M in total). Moreover, LLMut requires a moderate amount of input and output tokens, which cost 126 input tokens and 58-63 output tokens to generate one mutant. In total, LLMut needs 4.3M input tokens and 1.6M-2.1M output tokens, which are also moderate. Since LLMut asks the models to generate more mutants within a single conversation and wraps the output mutants directly with JSON tags, the average cost is significantly lower than BugFarm and LLMorpheus. These results demonstrate that various prompt engineering strategies significantly impact token costs.
3.4. RQ4: Categorizations of Non-Compilable Mutants
To answer this RQ, we first identify the error types of non-compilable mutants and then analyze the surrounding code contexts that are error-prone in generating them.
@@ 3.4.1. Categorizations of compilation error types
Non-compilable mutants require a compilation step to filter out, which results in wasted computational resources. As mentioned in Section 3.2, LLMs generate many non-compilable mutants. This RQ analyzes the types of errors and potential root causes of these non-compilable mutants. Following the setting of the previous steps, we first sample 384 non-compilation mutants from the outputs of GPT models, ensuring the confidence level is 95% and the margin of error is 5% (Linares-Vásquez et al., 2017; Tufano et al., 2019; Macklon et al., 2023; de Mello et al., 2015). From the manual analysis of these non-compilable mutants, we identified nine distinct error types, shown as follows.
@@
Incorrect Location: This type of error is due to a mutant being applied to an incorrect code location.
Incorrect Exceptions: This type of error is due to throwing incorrect exceptions.
@@
3.4.2. Categorizations of code context of non-compilable mutants
@@
To analyze which types of code are prone to causing LLMs to generate non-compilable mutants, we examined the code locations of all non-compilable mutants generated by all nine mutant generation approaches in Table 4, as shown in Figure 3. For most approaches (except Major and BugFarm), the code locations with MemberReference and MethodInvocation are the top-2 most-prevent AST node types. For example, for GPT and DeepSeek models, about 40% of non-compilable mutants occur at the location with MemberReference or MethodInvocation. This is potentially caused by the inherent complexity of these operations, which often involve multiple dependencies and references. If any required method or member is missing or incorrectly specified, it can easily lead to non-compileable mutants. The errors highlight a need for better context-aware mutant generation, ensuring that method calls and member references align with the intended program structure. Additionally, we inspect the deletion mutants rejected by the compiler and find that they account for 7.48% (GPT-3.5), 1.69% (GPT-4o), 0.41% (GPT-4o-Mini), 3.4% (DeepSeek-V3-671b), 1.11% (DeepSeek-Coder-V2-236b), 3.4% (BugFarm), 0.8% (LLMorpheus), 0.02% (StarChat), 0.14% (CodeLlama), 39.99% (LEAM), 0.11% (BERT), and 16.61% (Major) of all their non-compilable mutants, respectively. Thus, for LLMs, deletion is not the major reason for non-compilation.
3.5. RQ5: Distributions of Surviving Mutants
@@
As shown in Table 4, the mutation scores of rule-based approaches (i.e., PIT and Major) are lower than those of LLMs, indicating their mutants are harder to detect. Therefore, we intend to analyze the reasons for surviving mutants. To address this RQ, we adopt a setting similar to the analysis of equivalent mutants. Specifically, for each approach, we randomly sample its surviving mutants and manually classify them. We employ the categories of surviving mutant causes defined in the study by Just et al. (Just et al., 2014a) as follows.
@@
(1) Not covered, i.e., the mutant is located in code never executed by the test suite. We refer to the mutant set of this category as .
@@
(2) Covered but no state infection, i.e., the mutant is executed but does not propagate any state change. We refer to the mutant set of this category as .
@@
(3) State infection without detection, i.e., the mutant alters the program state but any test oracle does not observe the deviation. We refer to the mutant set of this category as .
@@
These categories are mutually exclusive, with the following relationships: , , and . This classification enables us to analyze the underlying reasons for surviving mutants and further evaluate the extent to which different approaches contribute to strengthening the test suite. Given a surviving mutant, we first determine whether it is equivalent; if so, it is discarded. For non-equivalent mutants, we then check whether they are uncovered. If a mutant is covered, we further examine whether it alters the program state, i.e., whether any memory values are modified. Finally, the remaining surviving mutants that are covered but do not lead to detectable differences are classified as not detected. This stepwise process ensures that every surviving mutant is uniquely assigned to one category (equivalent, uncovered, not injecting state, or not detected).
@@
As shown in Table 13, the distributions of surviving mutants differ significantly across approaches, where the columns NC%, NI%, and ND% represent the proportions of , , and within all non-equivalent mutants, respectively.
From the table, we can make the following observations. First, BugFarm and LLMut exhibit a substantially higher proportion of (i.e., not covered) mutants, while LLMopheus shows a higher proportion of (i.e., not injected) mutants. For instance, among the surviving mutants generated by BugFarm (GPT-4o) and BugFarm (DS-671b), 79.52% and 72.94% fall into the category, respectively. In contrast, the traditional approaches PIT, Major, and Bert exhibit much lower proportions of 46.24%, 45.65%, and 46.25%, respectively. This suggests that LLM-based approaches either tend to generate mutants located in previously untested code regions or, in the case of LLMorpheus, often produce mutations that do not alter the program state. This characteristic highlights the potential of LLMs in guiding test suite augmentation by directing attention to untested program areas. Notably, a recent study guides LLMs to generate unit test cases to kill surviving mutants (Harman et al., 2025; Wang et al., 2025b), indicating the usefulness of these mutants for other applications. Second, traditional approaches, especially rule-based methods such as PIT and Major, produce a larger fraction of (i.e., not injected) and (i.e., not detected) mutants compared to LLM-based methods. For instance, PIT and Major exhibit proportions of 29.03% and 32.61%, respectively, whereas both LLMut and BugFarm across different models remain below 20%. These types of surviving mutants require not only generating inputs that execute the mutated code but also propagating and detecting subtle semantic differences, which is inherently more challenging than increasing coverage. Therefore, the overall implication is that LLM-generated surviving mutants are more directly beneficial for strengthening unit tests, as they emphasize extending coverage, while rule-based approaches tend to stress the harder task of state propagation and detection.